Due to the COVID-19, the price of hand-sanitizer increases from 150 taka to 250 taka. In response, the

quantity demanded declines from 20 units to 15 units.

(a) Calculate the price elasticity of demand. (5)

(b) Explain if the demand is elastic or inelastic. Based on the elasticity, what change (increase,

decrease, or keep it the same) in price the seller can do to maximize his total revenue? Explain your

answer.


1
Expert's answer
2021-04-19T18:50:26-0400
"E=\frac{q_2-q_1}{p_2-p_1}\times\frac {p_2+p_1}{q_2+q_1}=-1.75"

Elastic demand. Buyers are sensitive to price changes. Demand in percentage terms changes more than price, that is, an increase in revenue is characteristic of elastic demand.